OpenStudy (anonymous):

Given a point and the slope, write an equation of the line in point-slope form. m = -4/5 (-4, 9) how do you solve this?

OpenStudy (saifoo.khan):

point slope form = y=mx+b

OpenStudy (saifoo.khan):

there's a shortcut, use: y - y1 = m(x-x1)\[ y - 9 = -\frac45(x-(-4)) \]

OpenStudy (anonymous):

y-9=-4/5(x+4) this is the answer i got

OpenStudy (saifoo.khan):

Now solve this for y

OpenStudy (anonymous):

having trouble. :/